# A Theoretical Framework for Grokking: Interpolation followed by Riemannian Norm Minimisation

### v1

@@ 5 Examples and experiments
Let $F(w)=\|Xw-y\|_{2}^{2}$ with $X\in\mathbb{R}^{n\times d}$, and assume that $\min F=0$. The problem is convex and the set of critical points is the affine subspace =$\mathcal{M}=\{w\,:\,Xw=y\}$; Assumption 2 is satisfied globally.
It is well known that unregularised gradient flow $w^{\mathrm{GF}}$ converges to $\mathcal{P}_{\mathcal{M}}(w_{0})$, the projection of the initial point on $\mathcal{M}$. Then, since $\mathcal{M}$ is convex, the Riemannian flow on $\mathcal{M}$ necessarily converges to the minimal $\ell_{2}$ norm solution $w^{\star}=X^{+}y$, where $X^{+}$ denotes the pseudo-inverse. Those two points are different (unless $w_{0}=0$), which leads to grokking, as $w^{\star}$ is expected to have better generalization properties than $\mathcal{P}_{\mathcal{M}}(w_{0})$ (Bartlett et al., 2020). In this setting, the trajectories of $w^{\lambda}$ can be computed explicitly to illustrate the two-timescale dynamics; see Appendix C.
@@
More precisely, we analyse the extreme overparameterised setting when $r=m+n$. In Appendix C, we show that the set $\mathcal{M}^{\star}$ of stationary points of $F$ which are nonsingular matrices forms a manifold. Provided that unregularised gradient flow converges to a nonsingular point, we can apply our results locally. These results state that, in the second, slow phase of the dynamics, the trajectories minimise $\|U\|_{F}^{2}+\|V\|_{F}^{2}$ on $\mathcal{M}^{\star}$. Recall that for a given matrix $M\in\mathbb{R}^{n\times m}$, we have
@@
where $\|M\|_{*}$ is the nuclear norm (Srebro et al., 2004, Lemma 1). Since minimising the nuclear norm promotes low-rank solutions, this indicates a drift toward low-rank matrices during the slow phase of the dynamics. In Appendix C, we study the more general class of matrix sensing problems and discuss an important technical subtlety: the set $\mathcal{M}^{\star}$ forms a manifold only after excluding degenerate points. Handling those singularities is highly non-trivial and remains an open direction for future work.
Figure 2 below empirically confirms this grokking for matrix completion. We here randomly generate a rank $3$ ground truth matrix $M^{\star}\in\mathbb{R}^{20\times 20}$, with non-zero singular values $\sigma^{\star}_{1},\sigma^{\star}_{2},\sigma^{\star}_{3}$. We randomly sample 50% of the entries to define the observed entries $\Omega$. We then perform gradient descent with weight decay parameter $\lambda=10^{-3}$ and stepsize $\gamma=10^{-2}$ on the loss $F(U,V)$ defined in Equation 8 and where the weights $U,V\in\mathbb{R}^{20\times 10}$ are initialised with i.i.d. standard Gaussian entries. We then track the training loss, the unmasked test loss $\|M^{\star}-UV^{\top}\|_{F}$, the weight norms $\|w\|^{2}=\|U\|_{F}^{2}+\|V\|_{F}^{2}$, and singular values of the reconstruction matrix $UV^{\top}$. Additional experimental details can be found in Section D.1.
@@
Although our theoretical framework does not allow for non-smooth architectures such as ReLU networks, we illustrate in Figure 3 that similar grokking dynamics can be observed in this case. We train a two-layer ReLU network of the form $f_{w}(x)=\sum_{j=1}^{m}u_{j}\,\mathrm{ReLU}(v_{j}x+b_{j}),$ with weights $w=(u,v,b)$ where the outer layer is $u\in\mathbb{R}^{m}$, the inner weight $v\in\mathbb{R}^{m}$ and bias $b\in\mathbb{R}^{m}$. The teacher function $f$ is a sum of $3$ ReLUs and is represented in dotted light blue in Figure 3. We generate a training dataset of $n=10$ points by sampling $x_{i}$ uniformly in $[-2,2]$ and computing $y_{i}=f(x_{i})$. These training points are shown as black crosses in Figure 3. We train the student network with $m=100$ by minimising the squared loss $F(w)=\tfrac{1}{2n}\sum_{i=1}^{n}\left(f_{w}(x_{i})-y_{i}\right)^{2}$ using gradient descent with weight decay $\lambda=10^{-3}$ for $T=10^{6}$ iterations and small step size. The initial weights are independently sampled from a Gaussian of variance $4$. At each iteration, we record the train loss, the $\ell_{2}$-norm of the weights, as well as the test loss over a fixed test dataset (plotted Figure 3, left).
Explaining the observed grokking phenomenon. At time $t_{1}=0$, the weights are randomly initialised and the training loss is high. By $t_{2}=1$, the training loss has dropped to nearly zero, and the iterates closely approximate the solution that would be obtained by unregularised gradient flow, this solution does not have a low norm and generalises poorly. Subsequently, around time $t=1/\lambda$, the weight norms begin to decrease, and by $t_{3}\approx 10^{5}$, they have drifted to a zero training loss solution which has a much lower $\ell_{2}$-norm and which generalises much better. Such solutions are believed to have a small number of "kinks" (Savarese et al., 2019; Parhi and Nowak, 2021; Boursier and Flammarion, 2023), as observed in Figure 3 (far right plot).
@@
We also study—both as an application of Theorem 1 and numerically—the architecture of diagonal neural networks in Appendices C and D.2, which serve as a toy problem for neural network training dynamics. In that case grokking promotes sparse estimators.
@@ 6 Conclusion
Large initialisations (NTK regime) are known to yield poor generalisation (Chizat et al., 2019; Liu et al., 2022c), while small initialisations (rich regime) can lead to slow convergence or convergence to suboptimal solutions for the training loss (Boursier and Flammarion, 2024a, b). Grokking may offer a desirable compromise, achieving fast convergence to an interpolating solution while retaining strong generalisation.
Lastly, our analysis focuses on regression settings with bounded dynamics. In classification tasks, by contrast, the stationary manifold lies “at infinity” once interpolation is achieved. Extending our approach to such settings remains an open and promising direction for future work, likely requiring techniques tailored to classification losses.
@@ A.1 Proof of Proposition 1
We first need to restrict the dynamics of $(w^{\lambda}(t))$ to some compact of $\mathbb{R}^{d}$. Without loss of generality we can inflate the compact in Assumption 1, so that we can assume there is some compact $K$ of $\mathbb{R}^{d}$ such that for all $t\geq 0$, $B(w^{\mathrm{GF}}(t),1)\subset K$, where $B(w^{\mathrm{GF}}(t),1)$ is the ball of radius $1$ centered at $w^{\mathrm{GF}}(t)$. We also define for any $\lambda>0$, $T_{\lambda}=\inf\{t\in\mathbb{R}_{+}\mid w^{\lambda}(t)\not\in K\}$.
@@
Given the definition of $T_{\lambda}$, Equation 9 and the fact that $\bigcup_{t\in\mathbb{R}_{+}}B(w^{\mathrm{GF}}(t),1)\subset K$, this then implies that for any $\lambda\leq\lambda^{\star}$, $T_{\lambda}>T$. In particular, for any $\lambda\leq\lambda^{\star}$, Equation 9 becomes
@@
Proposition 1 then directly follows. ∎
@@ A.2 Junction between fast and slow dynamics
In the limit $\lambda\to 0$, the whole fast dynamics described by Proposition 1 is crushed into the $t=0$ point of the slow timescale. While Sections 4.1 and 4.2 respectively provide descriptions of the fast and slow timescale dynamics, one needs to control the junction of these two dynamics. This junction is made possible by Lemma 2 below, as well as a timeshift argument in the proof of Proposition 2.
@@
If Assumption 1 holds, there exists a function $t(\lambda)$ such that $\lim_{\lambda\to 0}t(\lambda)=0$ and
@@
Lemma 2 states that for a well chosen timepoint $t(\lambda)$, which is of order $-\lambda\ln(\lambda)$, the slow dynamics solution $\tilde{w}^{\lambda}$ will be close to the limit of the unregularised flow at that timepoint. This result will then be key in showing that $\lim_{\lambda\to 0}\tilde{w}^{\lambda}$ admits a right limit in 00, which is given by $\lim_{t\to\infty}w^{\mathrm{GF}}(t)$.
Note that this order $-\lambda\ln(\lambda)$ for $t(\lambda)$ is necessary: a smaller value of $t(\lambda)$ would not allow enough time for the flow to approach the limit of the unregularised gradient flow; while a larger value would correspond to a time where the regularised flow significantly drifted from the unregularised one.
This time $-\lambda\ln(\lambda)$ can indeed be seen, in the fast timescale, as the point where the dynamics transition from mimicking the unregularised flow, to the slow dynamics that minimises the regularisation term within some manifold.
@@
Equation 9 in the proof of Proposition 1 yields that for any $t\in[0,T_{\lambda}]$, $\left\|w^{\lambda}(t)-w^{GF}(t)\right\|\leq\lambda te^{ct}\sup_{w\in K}\|w\|$. We can then define $t(\lambda)=\frac{-\lambda\ln\lambda}{2c}>0$ and observe that for any $t\leq\min(T_{\lambda},\frac{t(\lambda)}{\lambda})$,
@@
Since $\sqrt{\lambda}\ln(\lambda)\underset{\lambda\to 0}{\longrightarrow}0$, we have for $\lambda$ small enough that the above term is smaller than $1$. In particular for $\lambda$ small enough, $T_{\lambda}\geq\frac{t(\lambda)}{\lambda}$. From there, Equation 9 yields at $\frac{t(\lambda)}{\lambda}$ for any small enough $\lambda>0$
@@
Now note that our choice of $t(\lambda)$ is such that both the first and second term in the last inequality converge to 00 — indeed, $t(\lambda)e^{c\frac{t(\lambda)}{\lambda}}=\frac{-\sqrt{\lambda}\ln(\lambda)}{2c}\underset{\lambda\to 0}{\longrightarrow}0$ and $\frac{t(\lambda)}{\lambda}\underset{\lambda\to 0}{\longrightarrow}+\infty$ — which concludes the proof. ∎
A.3 Proof of Proposition 2
The main element to prove Proposition 2 is Lemma 4 below. We first need to state another auxiliary lemma, given by Lemma 3 below, and proven in Section B.5.
@@ A.3 Proof of Proposition 2
Consider Assumptions 1 and 2. Let $(u^{\lambda})_{\lambda>0}$ be a family of solutions of the following ODE for any $\lambda>0$ and $t\geq 0$,
@@
Note that the existence of a neighbourhood $U$ in Lemma 3 is guaranteed by Lemma 1. We can now state our key lemma.
@@
Consider Assumptions 1 and 2. Let $(u^{\lambda})_{\lambda>0}$ be a family of solutions of the following ODE for any $\lambda>0$ and $t\geq 0$,
@@
First consider a neighbourhood $U$ of $\mathcal{M}$ such that $\Phi$ is $\mathcal{C}^{2}$ on $U$ and Lemma 3 holds. From there thanks to Lemma 3, we can assume that $\lambda>0$ is chosen small enough so that $u^{\lambda}(t)\in U$ for any $t\in\mathbb{R}_{+}$. We can now compute the time derivative of $\Phi(u^{\lambda}(t))$, using the chain rule for any $t\in\mathbb{R}_{+}$ and $\lambda$ small enough:
@@
Then using the fact that for any $w\in U$, $D\Phi(w)\cdot\nabla F(w)=0$ [Li et al., 2021, Lemma C.2],
@@
It now remains to show that $u^{\lambda}(t)\to\mathcal{M}$ as $\lambda\to 0$, to guarantee that $\Phi(u^{\lambda}(t))$ and $u^{\lambda}(t)$ have the same limit, which will be done using Lemma 3.
Thanks to Lemma 3, we can consider a family of neighbourhoods $(U_{\varepsilon})_{\varepsilon>0}$ of $\mathcal{M}$ and a function $\lambda:\mathbb{R}_{+}^{*}\to\mathbb{R}_{+}^{*}$ satisfying Lemma 3. As we can always take a smaller choice for any value $\lambda(\varepsilon)$, we can also choose the function $\lambda$ so that
@@
For the remaining of proof, define the function $H:w\mapsto-D\Phi_{w}\cdot w$ and take $\varepsilon(\lambda)=\inf\{\varepsilon>0\mid\lambda\leq\lambda(\varepsilon)\}$. We consider in the following $\lambda$ small enough so that $\varepsilon(\lambda)$ is defined and finite. Since $\lim_{\varepsilon\to 0}\lambda(\varepsilon)=0$, $\varepsilon(\lambda)>0$ for any $\lambda>0$. The function $\varepsilon$ is non-increasing, so it admits a limit at 00. Moreover for any $\delta>0$, $\varepsilon(\lambda(\delta))\leq\delta$ by definition, sot that $\lim_{\lambda\to 0}\varepsilon(\lambda)=0$.
Thanks to Lemma 3, $(u^{\lambda}(t))_{t\geq 0}$ is contained in $U_{\varepsilon(\lambda)}$. By monotonicity of $\|\tilde{w}^{\circ}\|_{2}$, the trajectory $(\tilde{w}^{\circ}(t))_{t\geq 0}$ is bounded. Moreover, the trajectory of $\tilde{w}^{\lambda}(t)$ is also bounded independently of $\lambda$ thanks to Lemma 7. We can thus consider a compact $K$ of $\mathbb{R}^{d}$ such that for any $t\geq 0$ and $\lambda>0$, $\tilde{w}^{\circ}(t)\in K$ and $\tilde{w}^{\lambda}(t)\in K$.
Recall that $\Phi$ is the identity function on $\mathcal{M}$ and is $\mathcal{C}^{2}$ on $U$. In consequence, we have [^2] This is a direct consequence of Taylor’s expansion taken at the euclidean projection of $w$ on $\mathcal{M}$ and the fact that $\lim_{\lambda\to 0}\varepsilon(\lambda)=0$. $\sup_{w\in K\cap U_{\varepsilon(\lambda)}}\|\Phi(w)-w\|_{2}\underset{\lambda\to 0}{\longrightarrow}0$.
Summing over Equation 10 yields for any $t\geq 0$:
@@
Since $\Phi$ is $\mathcal{C}^{2}$ on $U$, $H$ is $c$-Lipschitz on $K$, for some $c>0$. A comparison with $u$ then yields for any $0\leq t\leq T_{\lambda}$
@@
Similarly to the proof of Proposition 1, an integral form of Grönwall inequality yields for any $t\in[0,T_{\lambda})$
@@
Noting that the multiplicative term $\sup_{w\in K\cap U_{\varepsilon(\lambda)}}\|\Phi(w)-w\|+\|\Phi(u^{\lambda}(0))-u_{0}\|$ goes to 00 as $\lambda$ goes to 00 allows to conclude on the uniform convergence of $u^{\lambda}$ to $u$ on $[0,T]$. ∎
@@
Consider the shifted slow dynamics $\tilde{v}^{\lambda}$ for any $t\geq 0$ as $\tilde{v}^{\lambda}(t)=\tilde{w}^{\lambda}(t+t(\lambda))$ with $t(\lambda)$ given by Lemma 2. Using Lemma 2, $\tilde{v}^{\lambda}$ then follows the following ODE:
@@
We can then direct apply Lemma 4 above on $\tilde{v}^{\lambda}$, which yields that $\tilde{v}^{\lambda}$ converges uniformly on any interval of the form $[0,T]$ to $\tilde{w}^{\circ}$.
Proposition 2 is then obtained by observing that $\lim_{\lambda\to 0}t(\lambda)=0$, so that for any $\varepsilon>0$ and $\lambda$ small enough such that $t(\lambda)\leq\varepsilon$, it holds for any $t\in[\varepsilon,T]$
@@
The first term converges to 00 uniformly for $t\in[\varepsilon,T]$ by uniform convergence of $\tilde{v}^{\lambda}$ towards $\tilde{w}^{\circ}$; and the second term also goes uniformly to 00 by (uniform) continuity of $\tilde{w}^{\circ}$ on the considered interval. ∎
A.4 Proof of Proposition 3

### v2

@@ 5 Examples and experiments
Let $F(w)=\|Xw-y\|_{2}^{2}$ with $X\in\mathbb{R}^{n\times d}$, and assume that $\min F=0$. The problem is convex and the set of critical points is the affine subspace $\mathcal{M}=\{w\,:\,Xw=y\}$; Assumption˜2 is satisfied globally.
It is well known that unregularised gradient flow $w^{\mathrm{GF}}$ converges to $\mathcal{P}_{\mathcal{M}}(w_{0})$, the projection of the initial point on $\mathcal{M}$ (lemaire1996asymptotical; gunasekar2018characterizing). Then, since $\mathcal{M}$ is convex, the Riemannian flow on $\mathcal{M}$ necessarily converges to the minimal $\ell_{2}$ norm solution $w^{\star}=X^{+}y$, where $X^{+}$ denotes the pseudo-inverse. Those two points are different (unless $w_{0}=0$), which leads to grokking, as $w^{\star}$ is expected to have better generalization properties than $\mathcal{P}_{\mathcal{M}}(w_{0})$ (bartlett2020benign). In this setting, the trajectories of $w^{\lambda}$ can be computed explicitly to illustrate the two-timescale dynamics; see Appendix˜C.
@@
More precisely, we analyse the extreme overparameterised setting when $r=m+n$. In Appendix˜C, we show that the set $\mathcal{M}^{\star}$ of stationary points of $F$ which are nonsingular matrices forms a manifold. Provided that unregularised gradient flow converges to a nonsingular point, we can apply our results locally. These results state that, in the second, slow phase of the dynamics, the trajectories minimise $\|U\|_{F}^{2}+\|V\|_{F}^{2}$ on $\mathcal{M}^{\star}$. Recall that for a given matrix $M\in\mathbb{R}^{n\times m}$, we have
@@
where $\|M\|_{*}$ is the nuclear norm (Srebro2004MaxMargin, Lemma 1). Since minimising the nuclear norm promotes low-rank solutions, this indicates a drift toward low-rank matrices during the slow phase of the dynamics. In Appendix˜C, we study the more general class of matrix sensing problems and discuss an important technical subtlety: the set $\mathcal{M}^{\star}$ forms a manifold only after excluding degenerate points. Handling those singularities is highly non-trivial and remains an open direction for future work.
Figure˜2 below empirically confirms this grokking for matrix completion. We here randomly generate a rank $3$ ground truth matrix $M^{\star}\in\mathbb{R}^{20\times 20}$, with non-zero singular values $\sigma^{\star}_{1},\sigma^{\star}_{2},\sigma^{\star}_{3}$. We randomly sample 50% of the entries to define the observed entries $\Omega$. We then perform gradient descent with weight decay parameter $\lambda=10^{-3}$ and stepsize $\gamma=10^{-2}$ on the loss $F(U,V)$ defined in Equation˜8 and where the weights $U,V\in\mathbb{R}^{20\times 10}$ are initialised with i.i.d. standard Gaussian entries. We then track the training loss, the unmasked test loss $\|M^{\star}-UV^{\top}\|_{F}$, the weight norms $\|w\|^{2}=\|U\|_{F}^{2}+\|V\|_{F}^{2}$, and singular values of the reconstruction matrix $UV^{\top}$. Additional experimental details can be found in Section˜D.1.
@@
Although our theoretical framework does not allow for non-smooth architectures such as ReLU networks, we illustrate in Figure˜3 that similar grokking dynamics can be observed in this case. We train a two-layer ReLU network of the form $f_{w}(x)=\sum_{j=1}^{m}u_{j}\,\mathrm{ReLU}(v_{j}x+b_{j}),$ with weights $w=(u,v,b)$ where the outer layer is $u\in\mathbb{R}^{m}$, the inner weight $v\in\mathbb{R}^{m}$ and bias $b\in\mathbb{R}^{m}$. The teacher function $f$ is a sum of $3$ ReLUs and is represented in dotted light blue in Figure˜3. We generate a training dataset of $n=10$ points by sampling $x_{i}$ uniformly in [2,2][-2,2] and computing $y_{i}=f(x_{i})$. These training points are shown as black crosses in Figure˜3. We train the student network with $m=100$ by minimising the squared loss $F(w)=\tfrac{1}{2n}\sum_{i=1}^{n}\left(f_{w}(x_{i})-y_{i}\right)^{2}$ using gradient descent with weight decay $\lambda=10^{-3}$ for $T=10^{6}$ iterations and small step size. The initial weights are independently sampled from a Gaussian of variance $4$. At each iteration, we record the train loss, the $\ell_{2}$-norm of the weights, as well as the test loss over a fixed test dataset (plotted Figure˜3, left).
Explaining the observed grokking phenomenon. At time $t_{1}=0$, the weights are randomly initialised and the training loss is high. By $t_{2}=1$, the training loss has dropped to nearly zero, and the iterates closely approximate the solution that would be obtained by unregularised gradient flow, this solution does not have a low norm and generalises poorly. Subsequently, around time $t=1/\lambda$, the weight norms begin to decrease, and by $t_{3}\approx 10^{5}$, they have drifted to a zero training loss solution which has a much lower $\ell_{2}$-norm and which generalises much better. Such solutions are believed to have a small number of "kinks" (savarese2019infinite; parhi2021banach; boursier2023penalising), as observed in Figure˜3 (far right plot).
@@
We also study—both as an application of Theorem˜1 and numerically—the architecture of diagonal neural networks in Appendices˜C and D.2, which serve as a toy problem for neural network training dynamics. In that case grokking promotes sparse estimators.
@@ 6 Conclusion
Large initialisations (NTK regime) are known to yield poor generalisation (chizat2019lazy; liu2022omnigrok), while small initialisations (rich regime) can lead to slow convergence or convergence to suboptimal solutions for the training loss (boursier2024early; boursier2024simplicity). Grokking may offer a desirable compromise, achieving fast convergence to an interpolating solution while retaining strong generalisation.
Note that our analysis is derived in the asymptotic regime $\lambda\to 0$, since this allows for a tractable analysis. Extending the theory to a fixed $\lambda$ is considerably harder, that said, in Appendix˜E we offer a heuristic analysis regarding how small $\lambda$ needs to be for grokking to emerge. Also note that our analysis can easily be extended to other types of regularisations. In particular, we believe empirical observations reported for training with Sharpness-Aware Minimization (andriushchenko2022towards, p. 7) may also be interpreted through the lens of grokking, albeit driven by SAM-style regularisation rather than standard weight decay. Lastly, our analysis focuses on regression settings with bounded dynamics. In classification tasks, by contrast, the stationary manifold lies “at infinity” once interpolation is achieved. Extending our approach to such settings remains an open and promising direction for future work, likely requiring techniques tailored to classification losses.
@@ A.1 Proof of Proposition˜1
We first need to restrict the dynamics of $(w^{\lambda}(t))$ to some compact of $\mathbb{R}^{d}$. Without loss of generality we can inflate the compact in Assumption˜1, so that we can assume there is some compact $K$ of $\mathbb{R}^{d}$ such that for all $t\geq 0$, $B(w^{\mathrm{GF}}(t),1)\subset K$, where $B(w^{\mathrm{GF}}(t),1)$ is the ball of radius $1$ centered at $w^{\mathrm{GF}}(t)$. We also define for any $\lambda>0$, $T_{\lambda}=\inf\{t\in\mathbb{R}_{+}\mid w^{\lambda}(t)\not\in K\}$.
@@
Given the definition of $T_{\lambda}$, Equation˜9 and the fact that $\bigcup_{t\in\mathbb{R}_{+}}B(w^{\mathrm{GF}}(t),1)\subset K$, this then implies that for any $\lambda\leq\lambda^{\star}$, $T_{\lambda}>T$. In particular, for any $\lambda\leq\lambda^{\star}$, Equation˜9 becomes
@@
Proposition˜1 then directly follows. ∎
@@ A.2 Junction between fast and slow dynamics
In the limit $\lambda\to 0$, the whole fast dynamics described by Proposition˜1 is crushed into the $t=0$ point of the slow timescale. While Sections˜4.1 and 4.2 respectively provide descriptions of the fast and slow timescale dynamics, one needs to control the junction of these two dynamics. This junction is made possible by Lemma˜2 below, as well as a timeshift argument in the proof of Proposition˜2.
@@
If Assumption˜1 holds, there exists a function $t(\lambda)$ such that $\lim_{\lambda\to 0}t(\lambda)=0$ and
@@
Lemma˜2 states that for a well chosen timepoint $t(\lambda)$, which is of order $\lambda\ln(\lambda)$, the slow dynamics solution $\tilde{w}^{\lambda}$ will be close to the limit of the unregularised flow at that timepoint. This result will then be key in showing that $\lim_{\lambda\to 0}\tilde{w}^{\lambda}$ admits a right limit in 0, which is given by $\lim_{t\to\infty}w^{\mathrm{GF}}(t)$.
Note that this order $\lambda\ln(\lambda)$ for $t(\lambda)$ is necessary: a smaller value of $t(\lambda)$ would not allow enough time for the flow to approach the limit of the unregularised gradient flow; while a larger value would correspond to a time where the regularised flow significantly drifted from the unregularised one.
This time $\lambda\ln(\lambda)$ can indeed be seen, in the fast timescale, as the point where the dynamics transition from mimicking the unregularised flow, to the slow dynamics that minimises the regularisation term within some manifold.
@@
Equation˜9 in the proof of Proposition˜1 yields that for any $t\in[0,T_{\lambda}]$, $\left\|w^{\lambda}(t)-w^{GF}(t)\right\|\leq\lambda te^{ct}\sup_{w\in K}\|w\|$. We can then define $t(\lambda)=\frac{-\lambda\ln\lambda}{2c}>0$ and observe that for any $t\leq\min(T_{\lambda},\frac{t(\lambda)}{\lambda})$,
@@
Since $\sqrt{\lambda}\ln(\lambda)\underset{\lambda\to 0}{\longrightarrow}0$, we have for $\lambda$ small enough that the above term is smaller than $1$. In particular for $\lambda$ small enough, $T_{\lambda}\geq\frac{t(\lambda)}{\lambda}$. From there, Equation˜9 yields at $\frac{t(\lambda)}{\lambda}$ for any small enough $\lambda>0$
@@
Now note that our choice of $t(\lambda)$ is such that both the first and second term in the last inequality converge to 0 — indeed, $t(\lambda)e^{c\frac{t(\lambda)}{\lambda}}=\frac{-\sqrt{\lambda}\ln(\lambda)}{2c}\underset{\lambda\to 0}{\longrightarrow}0$ and $\frac{t(\lambda)}{\lambda}\underset{\lambda\to 0}{\longrightarrow}+\infty$ — which concludes the proof. ∎
A.3 Proof of Proposition˜2
The main element to prove Proposition˜2 is Lemma˜4 below. We first need to state another auxiliary lemma, given by Lemma˜3 below, and proven in Section˜B.5.
@@ A.3 Proof of Proposition˜2
Consider Assumptions˜1 and 2. Let $(u^{\lambda})_{\lambda>0}$ be a family of solutions of the following ODE for any $\lambda>0$ and $t\geq 0$,
@@
Note that the existence of a neighbourhood $U$ in Lemma˜3 is guaranteed by Lemma˜1. We can now state our key lemma.
@@
Consider Assumptions˜1 and 2. Let $(u^{\lambda})_{\lambda>0}$ be a family of solutions of the following ODE for any $\lambda>0$ and $t\geq 0$,
@@
First consider a neighbourhood $U$ of $\mathcal{M}$ such that $\Phi$ is $\mathcal{C}^{2}$ on $U$ and Lemma˜3 holds. From there thanks to Lemma˜3, we can assume that $\lambda>0$ is chosen small enough so that $u^{\lambda}(t)\in U$ for any $t\in\mathbb{R}_{+}$. We can now compute the time derivative of $\Phi(u^{\lambda}(t))$, using the chain rule for any $t\in\mathbb{R}_{+}$ and $\lambda$ small enough:
@@
Then using the fact that for any $w\in U$, $D\Phi(w)\cdot\nabla F(w)=0$ [li2021happens, Lemma C.2],
@@
It now remains to show that $u^{\lambda}(t)\to\mathcal{M}$ as $\lambda\to 0$, to guarantee that $\Phi(u^{\lambda}(t))$ and $u^{\lambda}(t)$ have the same limit, which will be done using Lemma˜3.
Thanks to Lemma˜3, we can consider a family of neighbourhoods $(U_{\varepsilon})_{\varepsilon>0}$ of $\mathcal{M}$ and a function $\lambda:\mathbb{R}_{+}^{*}\to\mathbb{R}_{+}^{*}$ satisfying Lemma˜3. As we can always take a smaller choice for any value $\lambda(\varepsilon)$, we can also choose the function $\lambda$ so that
@@
For the remaining of proof, define the function $H:w\mapsto-D\Phi_{w}\cdot w$ and take $\varepsilon(\lambda)=\inf\{\varepsilon>0\mid\lambda\leq\lambda(\varepsilon)\}$. We consider in the following $\lambda$ small enough so that $\varepsilon(\lambda)$ is defined and finite. Since $\lim_{\varepsilon\to 0}\lambda(\varepsilon)=0$, $\varepsilon(\lambda)>0$ for any $\lambda>0$. The function $\varepsilon$ is non-increasing, so it admits a limit at 0. Moreover for any $\delta>0$, $\varepsilon(\lambda(\delta))\leq\delta$ by definition, so that $\lim_{\lambda\to 0}\varepsilon(\lambda)=0$.
Thanks to Lemma˜3, $(u^{\lambda}(t))_{t\geq 0}$ is contained in $U_{\varepsilon(\lambda)}$. By monotonicity of $\|u\|_{2}$, the trajectory $(u(t))_{t\geq 0}$ is bounded. Moreover, the trajectory of $u^{\lambda}(t)$ is also bounded independently of $\lambda$ thanks to Lemma˜7. We can thus consider a compact $K$ of $\mathbb{R}^{d}$ such that for any $t\geq 0$ and $\lambda>0$, $u(t)\in K$ and $u(t)\in K$.
Recall that $\Phi$ is the identity function on $\mathcal{M}$ and is $\mathcal{C}^{2}$ on $U$. In consequence, we have [^3] This is a direct consequence of the fact that $\Phi$ is the identity on $\mathcal{M}$, locally Lipschitz and $\lim_{\lambda\to 0}\varepsilon(\lambda)=0$. $\sup_{w\in K\cap U_{\varepsilon(\lambda)}}\|\Phi(w)-w\|_{2}\underset{\lambda\to 0}{\longrightarrow}0$.
Summing over Equation˜10 yields for any $t\geq 0$:
@@
Since $\Phi$ is $\mathcal{C}^{2}$ on $U$, $H$ is $c$-Lipschitz on $K$, for some $c>0$. A comparison with $u$ then yields for any $t\geq 0$
@@
Similarly to the proof of Proposition˜1, an integral form of Grönwall inequality yields for any $t\geq 0$
@@
Noting that the multiplicative term $\sup_{w\in K\cap U_{\varepsilon(\lambda)}}\|\Phi(w)-w\|+\|\Phi(u^{\lambda}(0))-u_{0}\|$ goes to 0 as $\lambda$ goes to 0 allows to conclude on the uniform convergence of $u^{\lambda}$ to $u$ on $[0,T]$. ∎
@@
Consider the shifted slow dynamics $\tilde{v}^{\lambda}$ for any $t\geq 0$ as $\tilde{v}^{\lambda}(t)=\tilde{w}^{\lambda}(t+t(\lambda))$ with $t(\lambda)$ given by Lemma˜2. Using Lemma˜2, $\tilde{v}^{\lambda}$ then follows the following ODE:
@@
We can then direct apply Lemma˜4 above on $\tilde{v}^{\lambda}$, which yields that $\tilde{v}^{\lambda}$ converges uniformly on any interval of the form $[0,T]$ to $\tilde{w}^{\circ}$.
Proposition˜2 is then obtained by observing that $\lim_{\lambda\to 0}t(\lambda)=0$, so that for any $\varepsilon>0$ and $\lambda$ small enough such that $t(\lambda)\leq\varepsilon$, it holds for any $t\in[\varepsilon,T]$
@@
The first term converges to 0 uniformly for $t\in[\varepsilon,T]$ by uniform convergence of $\tilde{v}^{\lambda}$ towards $\tilde{w}^{\circ}$; and the second term also goes uniformly to 0 by (uniform) continuity of $\tilde{w}^{\circ}$ on the considered interval. ∎
A.4 Proof of Proposition˜3